### v3

# MLR-Bench: Evaluating AI Agents on Open-Ended Machine Learning Research

Abstract
Recent advancements in AI agents have demonstrated their growing potential to drive and support scientific discovery. In this work, we introduce MLR-Bench, a comprehensive benchmark for evaluating AI agents on open-ended machine learning research. MLR-Bench includes three key components: (1) 201 research tasks sourced from NeurIPS, ICLR, and ICML workshops covering diverse ML topics; (2) MLR-Judge, an automated evaluation framework combining LLM-based reviewers with carefully designed review rubrics to assess research quality; and (3) MLR-Agent, a modular agent scaffold capable of completing research tasks through four stages: idea generation, proposal formulation, experimentation, and paper writing. Our framework supports both stepwise assessment across these distinct research stages, and end-to-end evaluation of the final research paper. We then use MLR-Bench to evaluate six frontier LLMs and an advanced coding agent, finding that while LLMs are effective at generating coherent ideas and well-structured papers, current coding agents frequently (e.g., in 80% of the cases) produce fabricated or invalidated experimental results—posing a major barrier to scientific reliability. We validate MLR-Judge through human evaluation, showing high agreement with expert reviewers, supporting its potential as a scalable tool for research evaluation. We open-source MLR-Bench to help the community benchmark, diagnose, and improve AI research agents toward trustworthy and transparent scientific discovery.
1 Introduction
A hallmark of human intelligence is the ability to generate new knowledge through research and scientific discovery. From basic research to applied technological development, scientific progress has shaped human civilization. Reproducing this ability—to autonomously generate, test, and validate new knowledge—has long been viewed as one of the grand challenges for artificial intelligence.
Recent advances in large language models (LLMs) bring us closer to this vision. LLM-powered agents have demonstrated impressive capabilities in generating research ideas (chainofideas2024, ; baek-etal-2025-researchagent, ; si2025can, ; li2024learning, ), conducting experiments and analyses (huang2024mlagentbench, ; wang2024openhands, ; chan2025mlebench, ; yang2024sweagent, ; jimenez2024swebench, ; yang2024swebenchmultimodal, ; jing2025dsbench, ), drafting scientific articles (radensky2024let, ; wang2024autosurvey, ; xi2025omnithinkexpandingknowledgeboundaries, ; agarwal2025litllms, ) and automated review (weng2025cycleresearcher, ; jin-etal-2024-agentreview, ; thakkar2025llmfeedbackenhancereview, ; darcy2024margmultiagentreviewgeneration, ; ye2024yetrevealingrisksutilizing, ). These advances move us beyond automating isolated tasks, opening up the possibility of automating the entire scientific process, from idea conception to experimentation to dissemination.
However, realizing this vision raises an important question: how do we rigorously evaluate the quality of research produced by AI agents? While recent works have shown promising results (starace2025paperbenchevaluatingaisability, ; seo2025paper2code, ; lu2024ai, ; schmidgall2025agent, ; Yamada2025ai, ), the community still lacks a comprehensive benchmark to systematically assess AI agents’ ability to conduct open-ended scientific research, making progress toward autonomous scientific discovery difficult to measure and compare fairly between agents. Moreover, there is also a need for empirical analysis to identify key failure modes—such as hallucinated results, lack of novelty, or methodological flaws — to quantify current limitations and inform future progress in research agents.
Figure 1: An overview of the framework of MLR-Bench, consisting of both an end-to-end evaluation (left) and a stepwise evaluation (right), each of which uses LLM judges to automatically assess performance over 201 tasks. For end-to-end evaluation, we use the same model as backbone in idea generation, proposal generation and paper writing. For stepwise evaluation, various models are tested and compared within each step.
Toward this vision, we introduce MLR-Bench, a comprehensive benchmark designed for evaluating AI agents on open-ended machine learning research tasks. Specifically, we seek to answer the following research questions:
  • RQ1: How well can AI agents conduct open-ended machine learning research?3)
  • RQ2: How effectively can an LLM judge evaluate research, as measured by its agreement with human reviewers?4)
  • RQ3: What are the key factors that affect the quality of AI-generated research? (§5)
To systematically address these questions, we decompose the research process into four successive stages: (1) idea generation, (2) research proposal formulation, (3) experimental execution and analysis, and (4) paper writing. This decomposition allows us to assess both stepwise progress and the overall end-to-end research capability of AI agents.
MLR-Bench consists of three key components: (1) a collection of 201 real-world research tasks sourced from NeurIPS, ICLR, and ICML workshops over the past three years, covering a wide range of machine learning domains, including LLMs, trustworthy AI, ML systems, AI for science; (2) MLR-Judge, an automated evaluation pipeline consisting of LLM-based judge and structured review rubrics; and (3) MLR-Agent, a modular research agent scaffold capable of automatically completing these tasks by following the four defined research stages.
To explore RQ1, we use MLR-Bench to evaluate six state-of-the-art models—including the most recent o4-mini (o4mini2025, ), Gemini-2.5-pro-preview (gemini2025, ), and Qwen3-235B-A22B (qwenteam2025qwen3, )—as well as the advanced coding agents, Claude Code (claude2025, ) and Codex (codex2025, ). Our results show that while these models perform well in generating ideas and writing papers, they often fail to deliver innovative or scientifically reliable research, mostly due to coding agents producing invalidated experimental results.
To validate the reliability of MLR-Judge (addressing RQ2), we recruit 10 machine learning experts with top-tier conference review experience to independently review the AI-generated research outputs. We find that the level of agreement between an LLM judge and a human judge, when assessing the same output, is very close to the level of agreement between two human judges, demonstrating its effectiveness as a reliable evaluation tool. Finally, to address RQ3, we analyze the justifications provided by both MLR-Judge and human reviewers, identifying a recurring failure mode where agents tend to generate fabricated or unverified results after execution failures—revealing a gap between fluent output generation and true scientific rigor.
Our contributions include:
  • MLR-Bench: To our knowledge, the most comprehensive evaluation benchmark for AI research agents to date, featuring 201 open-ended ML research tasks, a human-aligned MLR-Judge for automated research quality assessment, and a modular MLR-Agent supporting both stepwise and end-to-end research execution.
  • Comprehensive Evaluation and Insights: We apply MLR-Bench to evaluate 6 frontier models, providing comparative insights into their strengths and limitations in scientific discovery. Also, we compare MLR-Agent with AI Scientist V2 on MLR-Bench, and find existing research agents exhibit critical failure modes, including the widespread issue of fabricated experimental results and hallucinated methodology, pointing to key challenges for future research.
2 MLR-Bench
In this section, we describe the overall flow of MLR-Bench. Fig.˜1 presents an overview of the framework of MLR-Bench, which consists of both an end-to-end evaluation pipeline to evaluate AI agents’ ability to complete end-to-end research, and a stepwise evaluation pipeline, which separately evaluates AI agents’ abilities on (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing.
Tasks. To measure the performance of AI agents across various machine learning topics, we collect 201 tasks from ICLR, ICML, and NeurIPS workshops over the past three years. These tasks are categorized into 9 core ML topics: LLMs and Vision Language Models (VLMs), AI for Science, ML Theory, Trustworthy AI, Computer Vision, ML Systems, Multimodality, Reinforcement Learning (RL), and other emerging topics. Fig.˜2 illustrates the distribution of our tasks across these ML topics. To curate our dataset, we first review all workshops from ICLR, ICML, and NeurIPS conferences over the past three years. We then filter out duplicate workshops and select those that maintain complete information and target a general audience. Finally, we extract workshop overviews and topics to formulate our tasks.
End-to-End Evaluation. To evaluate AI agents’ abilities to fully automate the research process, we give them each of these tasks as input and require them to produce the final completed paper as output, which we evaluate using our MLR-Judge (Section˜2.1). To facilitate software experimentation, agents are provided with an environment that includes file system access, a Python runtime for executing scripts, and internet connectivity.
Stepwise Evaluation. To assess AI agents’ research capabilities in a more fine-grained manner, we evaluate them over 4 steps, using MLR-Judge to assess each step: (1) Idea Generation: we provide the agent with a research task and require it to generate a research idea. (2) Proposal Generation: we provide the agent with a research task and idea, and require it to generate a detailed research proposal. (3) Experimentation: the agent is given access to an environment for running experiments (similar to in the end-to-end case), and iteratively writes code and runs experiments. (4) Paper Writing: agents are given a full set of experimental outputs (report, figures, and a log of commands run during experimentation), and required to produce the final paper. This task involves multimodal input and output, so we test multimodal agents in this stage.
Each step of the stepwise evaluation requires a set of input data for the agents. For example, in step (1), the Idea Generation agents require research tasks as input data; we use the same 201 tasks used in end-to-end evaluation. In step (2) the Proposal Generation agents require tasks and research ideas as input. To construct this dataset, for each of the 201 tasks, we randomly sample one idea (generated by any Idea Generation agent) from step (1) for this task, resulting in 201 (task, idea) pairs that serve as the input for step (2). We apply the same procedure in steps (3) and (4), each time randomly sampling from the outputs of the previous step for a task, except in step (3) we manually select a subset of 10 (task, idea, proposal) triples based on their suitability for evaluating the experimentation agents, to keep the experiments manageable. Details about these 10 selected tasks can be found in Appendix˜A.
Figure 2: The number of tasks grouped by our ML primary categories.
Table 1: Review dimensions for different research processes.
Dimension Ideation Proposal Coding Writing End-to-End
Consistency
Clarity
Novelty
Feasibility
Completeness
Soundness
Insightfulness
Significance
Overall
2.1 MLR-Judge
In this work, we develop an automated evaluation tool, MLR-Judge, which consists of multiple carefully designed rubrics and an LLM judge. Prior to designing these rubrics, we summarize the key review dimensions that people focus on when evaluating research outputs at each step: Consistency, Novelty, Clarity, Feasibility, Completeness, Soundness, Insightfulness, Significance and Overall. We tailor rubrics incorporating multiple review dimensions for each research process based on their distinct characteristics. Fig.˜2 illustrates the review dimensions for different research processes.
Additionally, MLR-Judge incorporates an LLM judge alongside our rubrics. In this work, we select Gemini-2.5-Pro-Preview and Claude-3.7-Sonnet as our judge models, which have demonstrated superior reasoning capabilities and multimodal functionality, enabling them to review papers containing figures. These two judge models independently evaluate the quality of research outputs, and the final evaluation results are averaged from their individual assessments. At each research step, MLR-Judge evaluates the Consistency dimension by analyzing both the history information and the output of MLR-Agent. For the experimentation step, we examine the execution log of our coding agent to gain insights into the experimental process. Details of rubrics and prompts of MLR-Judge can be found in LABEL:app-subsec:judge-prompts.
2.2 MLR-Agent
In this work, we develop a simple and flexible agent scaffold, MLR-Agent, to evaluate the capabilities of different models in conducting open-ended research. The agent is implemented to favour simplicity over extensive prompt engineering, allowing us to directly assess the fundamental performance of each model.
MLR-Agent operates in two modes: a stepwise execution mode and an end-to-end execution mode. The stepwise execution mode enables us to evaluate each model’s performance in individual research steps, while the end-to-end execution mode allows us to assess the models’ capabilities in automating complete research projects. MLR-Agent’s automated research process follows the same research steps as illustrated in Fig.˜1, namely (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. MLR-Agent uses LLMs for steps (1) and (2), a coding agent (e.g., Claude Code) for step (3), and multimodal LLMs for step (4). Additionally, since most frontier models currently lack web search capabilities, they face challenges in accessing recent related work. Therefore, MLR-Agent uniformly employs GPT-4o-Search-Preview (searchgpt2024, ) to conduct a literature review between (1) Idea Generation and (2) Proposal Generation, providing reference materials for the agent. Details about the prompt used for each stage can be found in Section˜D.1.
3 How Well Can AI Agents Conduct Open-Ended Research?
This section describes empirical studies of how well AI agents perform on open-ended research. We use MLR-Agent as our agent scaffold to conduct research and evaluate a number of frontier models, as shown in Fig.˜3. In experimentation and analysis step, our coding agents execute on an Ubuntu 22.04 server with access to four NVIDIA RTX 3090 GPUs. We use MLR-Judge to evaluate the outcomes of different research processes. We report the average ratings of two judge models, as mentioned in Section˜2.1, and details of each rating can be found in Appendix˜C.
Table 2: Evaluated models (o4mini2025, ; claude2025, ; deepseekai2025deepseekr1incentivizingreasoningcapability, ; ministral2024, ; qwenteam2025qwen3, ; gemini2025, ; codex2025, ; geminicli2025, ) in different research stages.
Stage Evaluated Models
Ideation o4-mini, Claude-3.7-Sonnet, Deepseek-R1,
§3.1 Ministral-8B, Qwen3-235B-A22B,
Gemini-2.5-Pro-Preview
Proposal o4-mini, Claude-3.7-Sonnet, Deepseek-R1,
§3.2 Ministral-8B, Qwen3-235B-A22B,
Gemini-2.5-Pro-Preview
Coding Claude Code (Claude-3.7-Sonnet),
§3.3 Codex (o4-mini)
Writing o4-mini, Claude-3.7-Sonnet,
§3.4 Gemini-2.5-Pro-Preview
End-to-End o4-mini, Claude-3.7-Sonnet,
§3.5 Gemini-2.5-Pro-Preview
Figure 3: Scores of two LLM judge models across seven review dimensions on ten tasks.
3.1 Results of Idea Generation Evaluation
All models demonstrate strong performance in Consistency and Significance, while Feasibility and Novelty are more challenging. We test six frontier models on 201 tasks for idea generation and Table˜3 presents the evaluation results. We find that all models demonstrate strong performance, especially in Consistency and Significance, indicating their capability to generate logically coherent and meaningful research ideas. However, there exists a notable performance gap in Novelty and Feasibility, suggesting that generating innovative yet implementable ideas remains a significant challenge for current LLMs.
Model size may not be the sole determinant of idea generation quality. The results show that Deepseek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability, ) achieves the highest score in Consistency and Overall performance, while Gemini-2.5-Pro-Preview excels in Clarity and Feasibility. Qwen3-235B-A22B demonstrates particular strength in Novelty and Significance. Notably, even the smaller Ministral-8B (ministral2024, ) model achieves competitive performance in Feasibility, suggesting that model size may not be the sole determinant of idea generation quality.
Table 3: Evaluation results of six frontier LLMs averaged on 201 tasks in idea generation. Best and worst scores are highlighted in green and red background, respectively. o4-mini-high: o4-mini-2025-04-16 with reasoning_effort set to “high”.
Model Consistency Clarity Novelty Feasibility Significance Overall
Ministral-8B $8.99\pm 0.36$ $7.83\pm 0.50$ $6.66\pm 0.46$ $6.94\pm 0.67$ $8.36\pm 0.38$ $7.68\pm 0.40$
Deepseek-R1 $9.26\pm 0.29$ $8.25\pm 0.32$ $7.43\pm 0.40$ $6.93\pm 0.56$ $8.70\pm 0.31$ $8.11\pm 0.30$
Claude-3.7-Sonnet $9.13\pm 0.32$ $8.07\pm 0.37$ $7.39\pm 0.40$ $6.65\pm 0.58$ $8.69\pm 0.33$ $7.96\pm 0.33$
Qwen3-235B-A22B $9.20\pm 0.28$ $8.20\pm 0.32$ $7.62\pm 0.42$ $6.67\pm 0.52$ $8.73\pm 0.31$ $8.03\pm 0.29$
o4-mini-high $9.23\pm 0.28$ $8.23\pm 0.30$ $7.49\pm 0.41$ $7.01\pm 0.53$ $8.66\pm 0.33$ $8.11\pm 0.30$
Gemini-2.5-Pro-Preview $9.20\pm 0.31$ $8.27\pm 0.31$ $7.30\pm 0.37$ $7.11\pm 0.57$ $8.58\pm 0.35$ $8.08\pm 0.28$
3.2 Results of Proposal Generation Evaluation
All models excel at generating logically coherent research proposals with research significance, but novelty, soundness and feasibility are more challenging. We also evaluate six models on 201 tasks for proposal generation and results are presented in Table˜4. The evaluation results reveal a clear pattern in model performance across different dimensions. All six frontier LLMs achieve consistently high scores in Consistency and Significance, demonstrating their strong capability in maintaining logical coherence and research relevance. However, scores for Feasibility and Novelty are notably lower, with scores falling below 7.5 in most cases. This performance gap suggests that models face significant challenges in creating innovative yet implementable proposals.
Large reasoning models present greater capabilities in generating quality proposals. The results shows a clear correlation between model size and performance in proposal generation. The larger models (Gemini-2.5-Pro-Preview, o4-mini-high, Claude-3.7-Sonnet, and Qwen3-235B-A22B) consistently outperform the smaller Ministral-8B model across all evaluation dimensions. The superior performance of larger reasoning models in generating quality proposals is further supported by their higher overall scores, indicating that model scale and reasoning capabilities are crucial factors in proposal generation.
Table 4: Evaluation results of six frontier LLMs averaged on 201 tasks in proposal generation. Best and worst scores are highlighted in green and red background, respectively.
Model Consistency Clarity Novelty Soundness Feasibility Significance Overall
Ministral-8B $8.93\pm 0.22$ $7.65\pm 0.48$ $6.88\pm 0.67$ $7.03\pm 0.86$ $6.69\pm 0.80$ $8.53\pm 0.35$ $7.50\pm 0.48$
Deepseek-R1 $9.02\pm 0.19$ $8.20\pm 0.35$ $32\pm 0.64 7.7$$75\pm 0.57$ $6.96\pm 0.60$ $8.64\pm 0.36$ $8.02\pm 0.34$
Claude-3.7-Sonnet $9.05\pm 0.21$ $8.31\pm 0.31$ $7.48\pm 0.65$ $7.81\pm 0.56$ $6.75\pm 0.62$ $8.80\pm 0.36$ $8.04\pm 0.27$
Qwen3-235B-A22B $9.03\pm 0.21$ $8.17\pm 0.39$ $7.48\pm 0.61$ $7.66\pm 0.64$ $6.94\pm 0.64$ $8.69\pm 0.31$ $8.04\pm 0.32$
o4-mini-high $9.06\pm 0.17$ $8.34\pm 0.28$ $7.45\pm 0.56$ $7.90\pm 0.60$ $7.18\pm 0.67$ $8.68\pm 0.35$ $8.17\pm 0.28$
Gemini-2.5-Pro-Preview $9.10\pm 0.26$ $8.42\pm 0.29$ $7.55\pm 0.67$ $7.90\pm 0.57$ $6.95\pm 0.68$ $8.73\pm 0.40$ $8.16\pm 0.34$
3.3 Results of Experimentation Evaluation
Popular coding agents cannot yet produce solid experimental results. We evaluate the experimental execution by comparing Claude Code and Codex on 10 representative machine learning tasks selected from the 201 tasks. Our LLM judge assesses both the execution records and the final results. As presented in Table˜5, the overall ratings for both agents fall below the 6.0 acceptance threshold, indicating that current coding agents still struggle to generate robust experimental outcomes. Specifically, while Claude Code achieves respectable scores for Consistency, Completeness, and Novelty, it falters on Soundness, Insightfulness, and Significance. This weakness is detailed in Table˜5 and further highlighted in Fig.˜3, where the Gemini judge’s scores for its Soundness and Significance are below 3.0. This suggests that although Claude Code can devise and execute complete and novel experiments, the results often lack scientific robustness. In contrast, Codex performs well on Soundness but underperforms on Novelty, Insightfulness, and Significance. This suggests that while Codex is reliable in executing experiments, it lacks the capability to design novel ones.
Table 5: Evaluation results of Claude Code and Codex averaged on ten tasks in experiment execution. The scores are the average of two judge models.
Coding Agent Consistency Completeness Novelty Soundness Insightfulness Significance Overall
Claude Code $6.75\pm 1.00$ $6.00\pm 0.68$ $5.65\pm 0.82$ $4.75\pm 1.02$ $4.50\pm 0.97$ $4.70\pm 0.95$ $4.95\pm 0.82$
Codex $6.30\pm 1.46$ $5.05\pm 0.87$ $3.80\pm 1.19$ $6.15\pm 0.87$ $4.45\pm 1.28$ $3.40\pm 1.05$ $4.95\pm 1.02$
3.4 Results of Paper Writing Evaluation
Gemini-2.5-Pro-Preview presents superior performance in paper writing. We evaluate three models with multimodal capabilities on the same ten tasks as mentioned in Section˜3.3 in paper writing. The evaluation results presented in Table˜6 indicate that Gemini-2.5-Pro-Preview demonstrates stronger writing capabilities, while o4-mini-high shows relatively weaker performance. From manual observation, we believe this is partly due to Gemini-2.5-Pro-Preview’s detailed formal explanations, including algorithms, formulas, and derivations, along with its strong mathematical capabilities, while o4-mini-high performs worse due to its relatively terse style affecting its clarity. However, none of the models achieve high overall scores (above 7.0), which may be influenced by the weaker experimental results from the previous step. This demonstrates that experimental success is a critical determinant of overall research quality.
Table 6: Evaluation results of three models averaged on ten tasks in paper writing. Best and worst scores are highlighted in green and red background, respectively.
Model Consistency Clarity Completeness Soundness Overall
o4-mini-high $6.35\pm 1.13$ $7.25\pm 0.77$ $6.15\pm 0.77$ $5.05\pm 1.11$ $5.90\pm 0.97$
[…]
Claude-3.7-Sonnet $7.40\pm 0.95$ $7.80\pm 0.61$ $6.80\pm 0.64$ $5.85\pm 0.89$ $6.50\pm 0.81$
3.5 Results of End-to-End Evaluation
Current research agents struggle to generate robust scientific results.
We conduct an end-to-end evaluation to compare three reasoning models within our MLR-Agent scaffold and benchmark them against AI Scientist V2 (Yamada2025ai, ), with results summarized in Table˜7. Our findings reveal two key points. Firstly, when using the MLR-Agent scaffold, the overall ratings for all three models fall below the 6.0 acceptance threshold. Notably, all models receive their lowest scores on Soundness, which underscores their shared difficulty in producing scientifically sound results. Secondly, in a direct model comparison, o4-mini is significantly outperformed by Gemini and Claude, whose performances are comparable. This trend is inversely correlated with cost. When using Claude Code as the coding agent, o4-mini is the most affordable, followed by Gemini. Consequently, Gemini emerges as the most cost-effective model for end-to-end autonomous research, considering its strong performance and moderate cost.
Table 7: End-to-end evaluation results over the ten tasks. AI Scientist V2 is powered by o4-mini-high. We use o4-mini-medium in Codex and use Gemini-2.5-Pro-Preview-05-06 in this experiment. When using Claude Code as the coding agent, the costs of o4-mini-high, Gemini-2.5-Pro-Preview, and Claude-3.7-Sonnet are \$1.15, \$1.24, and \$2.40, respectively.
Model Clarity Novelty Soundness Significance Overall
AI Scientist V2
o4-mini-high $6.55\pm 0.94$ $6.70\pm 0.48$ $3.70\pm 1.29$ $4.85\pm 1.08$ $4.25\pm 1.25$
MLR-Agent
o4-mini-high + Codex $6.45\pm 0.90$ $5.65\pm 0.60$ $2.90\pm 0.57$ $3.80\pm 0.65$ $3.10\pm 0.60$
Gemini-2.5-Pro-Preview + Gemini CLI $8.30\pm 0.37$ $6.85\pm 0.34$ $4.15\pm 1.06$ $5.30\pm 0.91$ $4.60\pm 1.00$
Claude-3.7-Sonnet + Claude Code $7.75\pm 0.34$ $7.10\pm 0.43$ $4.05\pm 1.11$ $5.50\pm 1.14$ $4.70\pm 1.22$
AI Scientist V2 outperforms MLR-Agent, yet both share fundamental limitations.
We further compare MLR-Agent with AI Scientist V2 in Table˜7. Our comparison between MLR-Agent and AI Scientist V2, both using o4-mini as the backbone model, reveals that AI Scientist V2 consistently outperforms MLR-Agent across all review dimensions. However, neither agent meets the 6.0 acceptance threshold. A closer analysis of their detailed scores highlights three key findings:
  • Both agents excel at generating novel research outcome clearly. Both agents demonstrate strong creativity, reflected in high scores for Novelty. They can propose original research questions crucial for scientific discovery and articulate these ideas clearly, as shown by their strong performance in the Clarity dimension.
  • Agents struggle with technical soundness. A primary weakness is their inability to propose technically sound methods. This is evident from their markedly low scores in Soundness, which consistently fall below the acceptance threshold. This suggests that while the agents can formulate creative ideas, their technical execution is frequently flawed.
  • The agents’ research demonstrates limited potential impact. This limitation is reflected in their consistently low scores for Significance, suggesting their work often fails to address important research questions. This perceived lack of contribution, potentially compounded by reproducibility challenges, severely restricts the research’s influence and uptake by the scientific community.
4 How Well Is MLR-Judge Aligned with Human Reviewers?
To evaluate whether our proposed LLM-based judge, MLR-Judge, can serve as a reliable proxy for human reviewers, we conduct a human evaluation study with ten domain experts who have served as reviewers for major machine learning conferences such as NeurIPS, ICLR, or ICML. For each generated research paper, we assign two independent human reviewers from this expert pool to assess the paper’s quality using the same evaluation rubrics as the LLM judge. Specifically, all reviewers—whether human or MLR-Judge—follow the same review criteria, covering five dimensions: Clarity, Novelty, Soundness, Significance, and Overall. The detailed evaluation rubrics are provided in LABEL:tab:end2end_eval_prompt and the process is introduced in LABEL:sec:human-study.
Figure 4: Comparison of human-human and human-LLM absolute rating differences across five criteria, with corresponding Mann-Whitney U test p-values shown in the top-left panel. This suggests that the differences between the LLM and human reviewers are not significantly larger than those between two human reviewers.
Agreement Score. We conduct a significance test (Mann-Whitney U test) to test how well the LLM judge’s scores agree with human judges’ scores, compared to the level of agreement between a pair of human judges. Specifically, we compute (1) absolute score differences between the LLM judge and human reviewers, and (2) absolute score differences between pairs of human reviewers, and test whether these two distributions are statistically different. As shown in Fig.˜4, none of the five evaluation criteria show statistically significant differences at the 0.05 confidence level. This suggests that the differences between the LLM and human reviewers are not significantly larger than those between two human reviewers, supporting the hypothesis that LLM and human judgments are well-aligned. Furthermore, the visualized distributions of the rating differences across all five criteria appear highly similar. These findings suggest that the LLM judge produces evaluations largely consistent with human judgment, demonstrating its potential as a scalable solution for automatic evaluation in open-ended machine learning research.
5 What Are the Key Factors Affecting AI-Generated Research Quality?
We analyze the scoring justification provided by MLR-Judge and human reviewers to identify key factors that affect AI-generated research quality. Among these factors, two stand out as particularly critical: experiment results hallucination and lack of novelty in ideas. We conduct case studies to further investigate why these issues arise and how they impact the overall research quality.
Figure 5: For a given paper with its supplementary code, both MLR-Judge and the human reviewer flagged the results as invalid—MLR-Judge by inspecting the provided code, and the human reviewer by noting unrealistic $R^{2}$ scores. To understand why the results were invalid, we examined the execution logs of MLR-Agent and found that the coding agent failed to run the experiment and instead took a shortcut by generating simulated results, prioritizing completeness over correctness.
Experiment Results Hallucination. Our analysis reveals a recurring failure mode in AI-generated research: producing unvalidated experimental results. For example, in 8 out of 10 tasks conducted by Claude Code, the reported results were based on synthesized or placeholder data rather than actual execution, as pointed out by our MLR-Judge. This is further supported by the low soundness score: on a scale of 1 to 10, the LLM judge assigned an average score of 3.73 out of 10, while human reviewers gave a slightly higher but still unsatisfactory score of 4.42 out of 10.
Both human reviewers and MLR-Judge identified these issues. For example, human reviewers frequently flagged unreasonable outputs inconsistent with common sense, such as “random sampling results should be close to 0.5, but the paper shows 0.65, which seems fabricated”. Unlike human reviewers, who primarily focused on reading the paper content, MLR-Judge leveraged access to the supplementary code and detected even more hallucination cases by inspecting the execution logs and code traces of the coding agent, consistently assigned low scores on Soundness, Insightfulness, Significance, and Overall dimensions, as illustrated in Figure 3.
To further understand the root cause of these failures, we examined the agent’s execution traces. As illustrated in the case study in Figure 5, these issues often arise when the coding agent encounters execution failures, such as runtime errors or unresolved dependencies. Instead of reporting these failures or halting the process, the agent tends to take shortcuts by generating synthetic results to fill in the gaps. Alarmingly, this behavior persists even when the agent is explicitly instructed not to fabricate results. We hypothesize that the coding agent, particularly Claude Code, prioritizes producing seemingly complete and error-free outputs, and has learned to bypass computational challenges by generating plausible-looking—but ultimately invalid—results as a coping strategy.
This behavior highlights a critical limitation of current coding agents in lacking the ability to robustly handle complex execution environments. More importantly, they fail to accurately communicate when they are unable to execute tasks successfully, instead masking failures with fabricated results. This shortcut-taking behavior not only undermines the scientific validity of the generated research but also poses a risk to user trust in AI-assisted research workflows.
Lack of Novelty.
Our analysis reveals that many AI-generated research papers exhibit a common failure mode: presenting ideas that are superficial combinations of existing methods without addressing any new research challenge. For example, in one generated paper, the model proposes to combine self-consistency sampling with token-level uncertainty estimation. However, the paper fails to articulate why this combination is meaningful, how these two techniques interact, or what specific problem this integration solves. Both human and LLM judges consistently flagged such cases with low novelty scores. Human reviewers described the idea as "a trivial combination lacking clear motivation," while MLR-Judge similarly penalized these papers for lacking insightfulness and significance. Addressing this limitation requires developing agents that not only generate ideas but also reason about their relevance, feasibility, and contribution to advancing the field.
6 Related Work
Benchmarks for Autonomous Research Agents.
Recent works (radensky2024let, ; llmwritingsurvey2024, ; zhao2025sciarena, ) have explored using LLMs for various stages of scientific research, including idea generation (si2025can, ; ideabench2025, ; chainofideas2024, ), experimentation (huang2024mlagentbench, ; chan2025mlebench, ; tian2024scicode, ; miserendino2025swelancerfrontierllmsearn, ), and writing (llmwritingsurvey2024, ; xi2025omnithinkexpandingknowledgeboundaries, ; wang2024autosurvey, ). Existing benchmarks typically target narrow tasks: the MLE benchmark (chan2025mlebench, ) focuses on engineering, the MLAgentBench huang2024mlagentbench on experimentation, and the PaperBench starace2025paperbenchevaluatingaisability on paper reproduction, each revealing gaps between AI agents and human researchers (chan2025mlebench, ; huang2024mlagentbench, ; starace2025paperbenchevaluatingaisability, ). RE-Bench rebench2025 pushes toward generalization on unseen tasks (rebench2025, ). Our proposed MLR-Bench complements these efforts by evaluating 201 open-ended research tasks from major Machine Learning conferences, covering the entire research pipeline. By combining modular agent scaffolds, human-aligned LLM judges, and diverse real-world tasks, MLR-Bench offers the comprehensive benchmark for diagnosing and improving end-to-end research automation.
LLMs as Reviewers for Scientific Research. LLMs have also been explored as automated reviewers. Prior works like ReviewerGPT reviewergpt2024 and OpenReviewer openreviewer2024 show that LLMs can perform specific reviewing sub-tasks, but tend to produce overconfident and unreliable overall judgments (reviewergpt2024, ; openreviewer2024, ). More structured efforts, such as PaperBench starace2025paperbenchevaluatingaisability , demonstrate that fine-tuned LLM judges can align well with human reviewers on benchmark tasks (starace2025paperbenchevaluatingaisability, ). Building on these insights, MLR-Bench introduces MLR-Judge, a rubric-based LLM judge (zheng2023judging, ; chen2024mllmjudge, ; zhuge2024agent, ) validated against expert human ratings. Our study shows that MLR-Judge can be a proxy of human reviewers, enabling scalable and reliable evaluation of large-scale AI-generated research.
7 Conclusion
In this work, we introduce MLR-Bench, a comprehensive framework for evaluating AI research agents on open-ended machine learning research. MLR-Bench consists of two key components: (1) a collection of 201 real-world research tasks covering diverse ML topics, and (2) MLR-Judge, an LLM-based evaluation framework with structured review rubrics for automated assessment of research quality. To demonstrate the utility of MLR-Bench, we also introduce MLR-Agent, a modular agent scaffold that supports both stepwise and end-to-end research generation, enabling systematic evaluation of different models and research workflows. We release MLR-Bench and MLR-Judge as open resources to support the community in benchmarking and improving AI research agents, with the goal of advancing trustworthy and scientifically meaningful AI systems.
Limitation and Future Work. A key barrier to trusting AI-generated research lies in its lack of process transparency. Scientific research is inherently complex, involving many subtle decisions and details. Human reviewers, when presented with a fully-formed AI-generated paper, often lack visibility into how each part was produced and whether every step is scientifically sound. This makes it difficult to build trust in AI research agents. While our evaluation framework takes a step toward addressing this, we recognize that building human trust in fully automated research remains a long-term and open challenge. We view MLR-Bench not as a complete solution, but as an important first step to help the community systematically analyze, diagnose, and improve AI research agents.
Beyond evaluation, we believe that MLR-Bench and MLR-Judge have the potential to serve as valuable feedback signals for improving research agent training. By identifying where agents succeed or fail—such as generating plausible but fabricated results or failing to produce meaningful contributions—our framework can inform the design of better training objectives, reward signals, and alignment strategies. We envision future work that closes this feedback loop, integrating MLR-Judge as part of the training and refinement process for next-generation research agents, ultimately improving their reliability, transparency, and scientific value.
Acknowledgement
This research is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 2 (FY2025) (Grant MOE-T2EP20124-0009).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract has illustrated that this work aims to provide a benchmark for evaluating AI agents on open-ended machine learning research. We collect 201 tasks and test several frontier models in our benchmark.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Limitations section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work focuses on empirical studies and there are no theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We illustrated our main experiments in section 3.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We open-sourced our codebase and dataset at https://github.com/chchenhui/mlrbench and https://huggingface.co/datasets/chchenhui/mlrbench-tasks.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: MLR-Bench section illustrates our framework. More details can be found in Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: See section 4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See section 3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have checked the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Sure, we discuss our broader impact in Limitation and Conclusion.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We discuss these in Ethical Considerations.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the original paper or technical report that produced the code package or dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have uploaded our dataset to huggingface.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We have human evaluation and details can be found in this paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: Yes.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Yes, we described the usage of LLMs in the main content.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
Appendix A Experimental Details
This section introduces the details of our experiments, including the ten tasks we selected in Sections˜3.3, 3.4 and 3.5. Our experiments are conducted on an Ubuntu 22.04 server with access to four NVIDIA RTX 3090 GPUs.
Description of 10 Selected Tasks
This section describes the ten selected tasks in experimentation, paper writing and end-to-end evaluation. Ten tasks are selected from the most recent ICLR 2025 workshops and most of them are related to Trustworthy AI.
Table 8: Description of ten selected tasks in experimentation, paper writing and end-to-end evaluation.
Task Name Topic Category
iclr2025_bi_align Bidirectional Human-AI Alignment Trustworthy AI
iclr2025_buildingtrust Building Trust in Language Models and Applications Trustworthy AI
iclr2025_data_problems Navigating and Addressing Data Problems for Foundation Models Trustworthy AI
iclr2025_dl4c Emergent Possibilities and Challenges in Deep Learning for Code LLM/VLM
iclr2025_mldpr The Future of Machine Learning Data Practices and Repositories Trustworthy AI
iclr2025_question Quantify Uncertainty and Hallucination in Foundation Models: LLM/VLM
The Next Frontier in Reliable AI
iclr2025_scope Scalable Optimization for Efficient and Adaptive Foundation Models Trustworthy AI
iclr2025_scsl Spurious Correlation and Shortcut Learning: Foundations and Solutions Trustworthy AI
iclr2025_verifai VerifAI: AI Verification in the Wild Trustworthy AI
iclr2025_wsl Neural Network Weights as a New Data Modality ML Theory
Appendix B Hallucination Analysis of Existing Research Agents
To further investigate the issues in the content generated by AI Scientist V2 and MLR-Agent, we conducted a hallucination analysis in this section. Specifically, we aim to examine how frequently these research agents (i.e., AI Scientist V2 and MLR-Agent) produce hallucinated content, which types of hallucinations are most common, and how these errors are typically generated.
We observe that AI-generated research papers can fail to be trustworthy in several ways: for example, when the proposed method fails to address the stated motivation, or when the conclusions do not logically follow from the methodology. Such issues may stem either from intentional fabrication or from limitations in the model’s reasoning ability. To focus on objectively verifiable errors, we propose four frequently observed, fact-based hallucination types:
  • Faked Experimental Results: The data, metrics, or experiments’ outcomes are fabricated or never actually performed.
  • Hallucinated Methodology: The technical approaches are proposed but are not implemented (e.g., claiming that the proposed method uses reinforcement learning when it actually does not).
  • Incorrect Citations: The references to academic papers are incorrect or cannot be found.
  • Mathematical Errors: Incorrect equations, flawed derivations, or improper applications of mathematical concepts.
Figure 6: Frequency of each hallucination type in AI Scientist V2 and MLR-Agent across 10 tasks.
Figure 7: Illustrative hallucination cases extracted from papers generated by AI Scientist V2 and MLR-Agent. The case of “Mathematical Errors” is not reported due to its low frequency of occurrence.
To efficiently analyze the hallucination content, we first leverage two advanced automated judges—Gemini (Gemini-2.5-Pro-Preview-06-05) and Claude (Claude-3-7-Sonnet-20250219) to identify whether each generated paper contains any of these four hallucination types. Then the judgments are verified by our human annotators. To reduce the annotation burden, human annotators are provided with two review files, one from the Gemini judge and one from the Claude judge, which contain AI-detected hallucination types and their corresponding evidence. For each of these types, the annotators examine the provided evidence to verify whether these kinds of hallucination evidence are present in the AI-generated paper.
Key Observations.
Fig.˜6 reports the frequency of hallucination types across 10 benchmark tasks and Fig.˜7 presents representative examples of each type extracted from papers generated by AI Scientist V2 and MLR-Agent. We have two key observations.
Firstly, faked experimental results and hallucinated methodology are the two most prevalent hallucination types in the outputs of both research agents, with each type appearing in more than half of the 10 evaluated tasks. Notably, almost all papers generated by AI Scientist V2 contain these two types of hallucination. As illustrated in Fig.˜7, research agents may fabricate near-perfect results to achieve their research objectives, and in some cases, their implementations are not aligned with the proposed method and experimental setup. One possible reason is that the generated ideas lack feasibility, leaving the coding agents unable to utilize available resources to achieve the experimental objectives. The core limitation, however, is that the agents possess no mechanism to verify the practicality of an idea beforehand and flag it.
Secondly, nonexistent citations are a significant issue, appearing in 50% of the tasks completed by MLR-Agent. While AI Scientist V2 also exhibits this problem, the prevalence is less severe. We hypothesize that this discrepancy stems from limitations in MLR-Agent’s literature collection tool, which restricts its ability to conduct accurate searches for relevant literature on the web. Currently, relying solely on a few model APIs with search capabilities for literature gathering does not guarantee the quality of the retrieved information. Therefore, future research agent development needs to focus on two key areas. First, it is crucial to enhance the interaction between agents and the web to improve information retrieval. Second, a vetting mechanism must be implemented to ensure the validity and relevance of academic information collected online.
Appendix C Evaluation Results of Different Judge Models
This section shows the evaluation scores of different models judged by Gemini and Claude models. We report the results of MLR-Agent in idea generation (see Tables˜9 and 10), proposal generation (see Tables˜11 and 12), paper writing (see Tables˜13 and 14) and end-to-end evaluation (see Tables˜15 and 16). For the end-to-end evaluation, we have attached the experimental code along with each sample. In addition,  Table˜17 and  Table˜18 present the performance comparison of AI Scientist V2 and MLR-Agent.
Table 9: Evaluation results judged by Gemini-2.5-Pro-Preview-03-25 in idea generation.
Model Consistency Clarity Novelty Feasibility Significance Overall
Ministral-8B $9.22\pm 0.57$ $8.01\pm 0.89$ $6.57\pm 0.74$ $6.77\pm 1.01$ $8.61\pm 0.52$ $7.75\pm 0.63$
Deepseek-R1 $9.57\pm 0.52$ $8.61\pm 0.54$ $7.48\pm 0.56$ $6.92\pm 0.81$ $8.87\pm 0.36$ $8.29\pm 0.51$
Claude-3.7-Sonnet $9.34\pm 0.58$ $8.26\pm 0.65$ $7.42\pm 0.56$ $6.56\pm 0.87$ $8.80\pm 0.44$ $8.00\pm 0.59$
Qwen3-235B-A22B $9.45\pm 0.51$ $8.49\pm 0.57$ $7.66\pm 0.54$ $6.65\pm 0.78$ $8.83\pm 0.40$ $8.12\pm 0.50$
o4-mini-high $9.52\pm 0.50$ $8.55\pm 0.52$ $7.52\pm 0.58$ $6.96\pm 0.76$ $8.80\pm 0.41$ $8.29\pm 0.54$
Gemini-2.5-Pro-Preview $9.46\pm 0.57$ $8.59\pm 0.52$ $7.34\pm 0.56$ $7.00\pm 0.78$ $8.68\pm 0.49$ $8.21\pm 0.50$
Table 10: Evaluation results judged by Claude-3.7-Sonnet-20250219 in idea generation.
Model Consistency Clarity Novelty Feasibility Significance Overall
Ministral-8B $8.75\pm 0.43$ $7.64\pm 0.48$ $6.74\pm 0.55$ $7.10\pm 0.89$ $8.11\pm 0.54$ $7.61\pm 0.51$
Deepseek-R1 $8.94\pm 0.24$ $7.89\pm 0.33$ $7.37\pm 0.57$ $6.94\pm 0.78$ $8.52\pm 0.50$ $7.93\pm 0.28$
Claude-3.7-Sonnet $8.92\pm 0.27$ $7.88\pm 0.35$ $7.36\pm 0.56$ $6.73\pm 0.75$ $8.57\pm 0.49$ $7.91\pm 0.30$
Qwen3-235B-A22B $8.94\pm 0.26$ $7.91\pm 0.30$ $7.57\pm 0.64$ $6.69\pm 0.67$ $8.63\pm 0.48$ $7.93\pm 0.29$
o4-mini-high $8.94\pm 0.25$ $7.90\pm 0.30$ $7.46\pm 0.58$ $7.06\pm 0.73$ $8.52\pm 0.52$ $7.92\pm 0.27$
Gemini-2.5-Pro-Preview $8.93\pm 0.26$ $7.94\pm 0.33$ $7.26\pm 0.47$ $7.21\pm 0.82$ $8.47\pm 0.51$ $7.95\pm 0.27$
Table 11: Evaluation results judged by Gemini-2.5-Pro-Preview-03-25 in proposal generation.
Model Consistency Clarity Novelty Soundness Feasibility Significance Overall
Ministral-8B $8.94\pm 0.33$ $7.32\pm 0.90$ $6.72\pm 1.30$ $6.46\pm 1.13$ $6.55\pm 0.93$ $8.78\pm 0.48$ $7.21\pm 0.90$
Deepseek-R1 $9.04\pm 0.26$ $8.39\pm 0.58$ $7.44\pm 0.83$ $7.68\pm 0.72$ $7.05\pm 0.77$ $8.97\pm 0.24$ $8.04\pm 0.54$
Claude-3.7-Sonnet $9.09\pm 0.33$ $8.62\pm 0.52$ $7.56\pm 0.77$ $7.90\pm 0.55$ $6.77\pm 0.78$ $9.01\pm 0.20$ $8.08\pm 0.52$
Qwen3-235B-A22B $9.07\pm 0.30$ $8.33\pm 0.64$ $7.59\pm 0.71$ $7.66\pm 0.77$ $7.02\pm 0.84$ $9.00\pm 0.10$ $8.08\pm 0.50$
o4-mini-high $9.12\pm 0.33$ $8.68\pm 0.54$ $7.57\pm 0.64$ $8.03\pm 0.68$ $7.37\pm 0.86$ $8.99\pm 0.20$ $8.34\pm 0.55$
Gemini-2.5-Pro-Preview $9.21\pm 0.41$ $8.82\pm 0.42$ $7.68\pm 0.81$ $8.18\pm 0.62$ $7.15\pm 0.81$ $9.04\pm 0.31$ $8.32\pm 0.57$
Table 12: Evaluation results judged by Claude-3.7-Sonnet-20250219 in proposal generation.
Model Consistency Clarity Novelty Soundness Feasibility Significance Overall
Ministral-8B $8.92\pm 0.29$ $7.97\pm 0.30$ $7.03\pm 0.31$ $7.59\pm 0.59$ $6.82\pm 0.67$ $8.28\pm 0.50$ $7.78\pm 0.43$
Deepseek-R1 $8.99\pm 0.10$ $8.00\pm 0.10$ $7.20\pm 0.43$ $7.82\pm 0.39$ $6.86\pm 0.40$ $8.30\pm 0.46$ $7.99\pm 0.12$
Claude-3.7-Sonnet $9.00\pm 0.07$ $8.00\pm 0.07$ $40\pm 0.51 7.7$$72\pm 0.57$ $6.73\pm 0.45$ $8.58\pm 0.49$ $8.00\pm 0.00$
Qwen3-235B-A22B $8.99\pm 0.10$ $8.00\pm 0.10$ $7.36\pm 0.50$ $7.66\pm 0.49$ $6.85\pm 0.42$ $8.37\pm 0.48$ $7.99\pm 0.12$
o4-mini-high $9.00\pm 0.00$ $8.00\pm 0.00$ $33\pm 0.48 7.7$$77\pm 0.52$ $6.98\pm 0.47$ $8.37\pm 0.48$ $8.00\pm 0.00$
Gemini-2.5-Pro-Preview $8.99\pm 0.10$ $8.02\pm 0.14$ $7.42\pm 0.51$ $7.62\pm 0.52$ $6.75\pm 0.54$ $8.41\pm 0.49$ $7.99\pm 0.10$
Table 13: Evaluation results judged by Gemini-2.5-Pro-Preview-05-06 in paper writing.
Model Consistency Clarity Completeness Soundness Overall
o4-mini-high $5.10\pm 1.76$ $6.60\pm 1.43$ $5.30\pm 1.27$ $4.00\pm 1.73$ $4.80\pm 1.47$
Gemini-2.5-Pro-Preview $7.10\pm 1.81$ $7.80\pm 0.98$ $6.60\pm 1.43$ $5.60\pm 1.11$ $5.90\pm 1.51$
Claude-3.7-Sonnet $6.70\pm 1.79$ $7.10\pm 1.22$ $5.90\pm 1.14$ $5.00\pm 1.73$ $5.40\pm 1.62$
Table 14: Evaluation results judged by Claude-3.7-Sonnet-20250219 in paper writing.
Model Consistency Clarity Completeness Soundness Overall
o4-mini-high $7.60\pm 1.43$ $7.90\pm 0.83$ $7.00\pm 0.89$ $6.10\pm 1.37$ $7.00\pm 1.26$
Gemini-2.5-Pro-Preview $8.00\pm 1.00$ $8.30\pm 0.46$ $7.80\pm 0.60$ $6.50\pm 0.92$ $7.30\pm 0.90$
Claude-3.7-Sonnet $8.10\pm 0.70$ $8.50\pm 0.50$ $7.70\pm 0.64$ $6.70\pm 0.78$ $7.60\pm 0.66$
Table 15: End-to-end evaluation results judged by Gemini-2.5-Pro-Preview-05-06.
Model Clarity Novelty Soundness Significance Overall
o4-mini-high $7.30\pm 0.64$ $6.80\pm 0.60$ $2.00\pm 1.00$ $3.50\pm 0.92$ $2.20\pm 1.17$
Gemini-2.5-Pro-Preview $7.00\pm 0.89$ $6.80\pm 1.17$ $2.00\pm 1.55$ $3.20\pm 1.78$ $2.30\pm 1.42$
Claude-3.7-Sonnet $7.50\pm 0.50$ $7.20\pm 0.75$ $2.80\pm 1.99$ $4.40\pm 2.29$ $3.50\pm 2.29$
Table 16: End-to-end evaluation results judged by Claude-3.7-Sonnet-20250219.
Model Clarity Novelty Soundness Significance Overall
o4-mini-high $8.00\pm 0.00$ $7.00\pm 0.00$ $5.60\pm 0.66$ $6.60\pm 0.66$ $5.70\pm 0.78$
Gemini-2.5-Pro-Preview $7.80\pm 0.40$ $6.70\pm 0.46$ $4.70\pm 0.78$ $5.90\pm 0.54$ $5.20\pm 0.87$
Claude-3.7-Sonnet $8.00\pm 0.45$ $7.00\pm 0.45$ $5.30\pm 1.00$ $6.60\pm 0.80$ $5.90\pm 1.14$
Table 17: End-to-end performance comparison of AI Scientist V2 and MLR-Agent judged by Gemini-2.5-Pro-Preview-05-06. Each score is averaged across 10 tasks.
Review Dimensions
Agent Scaffold Clarity Novelty Soundness Significance Overall
AI Scientist V2 $6.50\pm 1.69$ $6.90\pm 0.83$ $00\pm 2.37 5.5$$50\pm 1.75$ $5.00\pm 2.19$
MLR-Agent $7.30\pm 0.64$ $6.80\pm 0.60$ $2.00\pm 1.00$ $3.50\pm 0.92$ $2.20\pm 1.17$
Table 18: End-to-end performance comparison of AI Scientist V2 and MLR-Agent judged by Claude-3.7-Sonnet-20250219. Each score is averaged across 10 tasks.
Review Dimensions
Agent Scaffold Clarity Novelty Soundness Significance Overall
AI Scientist V2 $6.60\pm 0.80$ $6.50\pm 0.50$ $2.40\pm 1.02$ $4.20\pm 1.25$ $3.50\pm 1.20$
MLR-Agent $8.00\pm 0.00$ $7.00\pm 0.00$ $5.60\pm 0.66$ $6.60\pm 0.66$ $5.70\pm 0.78$
Appendix D Prompts
In this section, we present the prompts used throughout the MLR-Bench pipeline. Fields highlighted in {blue} indicate parts that require user-specified input or content specified in another place. To make the process more intuitive, we also include example inputs and outputs where helpful. In cases where the content is lengthy, we use ellipses (“…”) to omit parts of the input or output for clarity.
D.1 Prompts for MLR-Agent
MLR-Agent is a simple and flexible scaffold designed to support open-ended research. It is implemented to favour simplicity over extensive prompt engineering, allowing us to assess each large language model’s fundamental performance directly.
The automated research process includes four stages: (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. Since most models lack web access, we insert a literature review step using GPT-4o-Search-Preview (searchgpt2024, ) between stages (1) and (2), providing relevant reference material to the agent.
Below are the prompts used for each stage:
  • LABEL:tab:idea_prompt presents the prompt used for idea generation.
  • LABEL:tab:lit_review_prompt presents the prompt for instructing the model to perform a literature review.
  • LABEL:tab:proposal_prompt presents the prompt used for generating the research proposal.
  • LABEL:tab:experiment_prompt presents the prompt used for conducting experiments and obtaining results.
  • LABEL:tab:writer_prompt presents the prompt used for writing the research paper based on previously generated outputs, including the idea, proposal, and experimental results.
Table 33: The Google Form interface used to collect review scores, following the same reviewer rubrics as MLR-Judge (LABEL:tab:end2end_eval_prompt).
Task # Workshop on Building Trust in Language Models and Applications
Input As Large Language Models (LLMs) are rapidly adopted across diverse industries, concerns around their trustworthiness, safety, and ethical implications increasingly motivate academic research, industrial development, and legal innovation…This workshop addresses the unique challenges posed by the deployment of LLMs, ranging from guardrails to explainability to regulation and beyond. The proposed workshop will bring together researchers and practitioners from academia and industry to explore cutting-edge solutions for improving the trustworthiness of LLMs and LLM-driven applications. The workshop will feature invited talks, a panel discussion, interactive breakout discussion sessions, and poster presentations, fostering rich dialogue and knowledge exchange. We aim to bridge the gap between foundational research and the practical challenges of deploying LLMs in trustworthy, use-centric systems.
## Workshop Scope
This workshop has a broad focus, including but not limited to:
1. Metrics, benchmarks, and evaluation of trustworthy LLMs
2. Improving reliability and truthfulness of LLMs
3. Explainability and interpretability of language model responses
7. Guardrails and regulations for LLMs
8. Error detection and correction
Prompt You are an excellent machine learning researcher. Please generate innovative and practical ideas based on a given task description.
Note that there might be a couple of research topics in the task description, and you should focus on one of them.
The idea should be no more than 200 words and include the following three sections:
1. Title: A concise and descriptive title for the research idea.
2. Motivation: A brief explanation of why this research is important and what problems it aims to solve.
3. Main Idea: A clear and detailed description of the proposed research idea, including the methodology, expected outcomes, and potential impact.
Please directly respond to the idea.
Here is the task: {Task}
Output # TrustPath: A Framework for Transparent Error Detection and Correction in LLMs
## Motivation
As Large Language Models (LLMs) are increasingly integrated into high-stakes applications, their tendency to hallucinate, present false information, or make errors poses significant risks. Current error detection methods often lack transparency, making it difficult for users to understand when and why an LLM might be incorrect. This research addresses the critical need for systems that can not only detect errors in LLM outputs but also explain these errors to users and suggest corrections in a transparent, understandable manner.
## Main Idea
TrustPath proposes a multi-layered approach to error detection and correction that prioritizes transparency. The framework employs three complementary mechanisms: 1) A self-verification module that prompts the LLM to evaluate its own uncertainty through generated alternative responses and explanations; 2) A factual consistency checker that retrieves information from trusted knowledge sources to verify claims; and 3) A human-in-the-loop feedback system that learns from user corrections. Uniquely, TrustPath presents detected errors alongside confidence scores, reasoning paths, and suggested alternatives through an intuitive visual interface that highlights potential problematic sections of text. This approach not only improves accuracy but builds user trust by making the error detection process visible and understandable, allowing users to make informed decisions about accepting or rejecting LLM outputs.
Table 33: The Google Form interface used to collect review scores, following the same reviewer rubrics as MLR-Judge (LABEL:tab:end2end_eval_prompt).
Prompt You are an excellent machine learning researcher!
Please help me do a literature review for a given idea.
The idea is based on a given task description.
The papers in the literature review should be extracted from arxiv and published between 2023 to 2025.
The literature review should include the following:
1. Related Papers: at least 10 academic papers most closely related to the current research idea, with a brief summary and the publication year of each one, organized logically.
2. Key Challenges: A discussion of the main challenges and limitations in the current research. List no more than five key challenges.
The paper should be in the format of:
1. Title: <title> (<arxiv_id>)
- Authors: <author1>, <author2>, …
- Summary: <summary>
- Year: <year>
Please directly respond to the literature review and do not include any additional comments.
Here is the idea: {Idea}
Here is the task: {Task}
Output 1. Title: Explainability for Large Language Models: A Survey (arXiv:2309.01029)
- Authors: Haiyan Zhao, Hanjie Chen, Fan Yang, Ninghao Liu, Huiqi Deng, Hengyi Cai, Shuaiqiang Wang, Dawei Yin, Mengnan Du
- Summary: This survey introduces a taxonomy of explainability techniques for Transformer-based language models, categorizing methods based on training paradigms and summarizing approaches for generating local and global explanations. It also discusses evaluation metrics and the use of explanations to debug models and improve performance.
- Year: 2023
10. Title: An Explainable AI Approach to Large Language Model Assisted Causal Model Auditing and Development (arXiv:2312.16211)
- Authors: Yanming Zhang, Brette Fitzgibbon, Dino Garofolo, Akshith Kota, Eric Papenhausen, Klaus Mueller
- Summary: This study proposes the use of LLMs as auditors for causal networks, presenting them with causal relationships to produce insights about edge directionality, confounders, and mediating variables. It envisions a system where LLMs, automated causal inference, and human analysts collaborate to derive comprehensive causal models.
- Year: 2023
Key Challenges:
1. Transparency and Interpretability: LLMs often operate as black boxes, making it difficult to understand their internal decision-making processes. This lack of transparency hinders trust and accountability in high-stakes applications.
5. User Feedback Mechanisms: Incorporating human-in-the-loop feedback systems to learn from user corrections and improve model performance requires designing interfaces and processes that are both effective and user-friendly.
Table 33: The Google Form interface used to collect review scores, following the same reviewer rubrics as MLR-Judge (LABEL:tab:end2end_eval_prompt).
Prompt You are an excellent machine learning researcher!
Please generate a detailed research proposal based on a given task description, a research idea and their literature review.
The proposal should be about 2000 words and include the following five sections:
1. Title: a concise and descriptive title for the research proposal.
2. Introduction: background, research objectives and significance.
3. Methodology: detailed and precise research design (including data collection, full algorithmic steps and/or mathematical formulas where appropriate, and full details about experimental design to validate the method, with evaluation metrics).
4. Expected Outcomes & Impact.
The proposal should be well-structured and clearly articulate the research plan. When writing mathematical formulas, you should use LaTeX syntax. For inline formulas, use single dollar signs, for example: \$xˆ2\$ to represent x squared. For block equations, use double dollar signs at the beginning and end, for example: \$\$xˆ2\$\$.
Please directly respond to the proposal.
Here is the task: {Task}
Here is the idea: {Idea}
Here is the literature review: {Related Work}
Table 33: The Google Form interface used to collect review scores, following the same reviewer rubrics as MLR-Judge (LABEL:tab:end2end_eval_prompt).
Prompt Based on task.md, idea.md, related_work.md and proposal.md in this folder, please design an experimental plan to test the hypothesis that the proposed method is effective. Then write code to implement the experimental plan, and run the experiments in a fully automated manner.
The experimental process should be fully automated and include the following steps:
1. Create a folder named ’claude_code’ inside this folder to save the code and results.
2. Write some Python scripts to run the experiment automatically. IMPORTANT: The script must be thoroughly debugged and tested until it can run successfully without any errors. The script should handle all necessary data processing, model training, and evaluation steps automatically. NOTE: If the environment has available GPUs, the script should utilize GPU acceleration where possible.
- The scripts should be able to run all baseline methods automatically. - The scripts should be able to save the results in a structured format (e.g., CSV or JSON) for easy analysis.
- The scripts should generate figures to visualize the results. Figures should be generated for the main results, including but not limited to:
- Training and validation loss curves
- Performance metrics (e.g., accuracy, F1 score) over time
- Comparison of the proposed method with baseline methods
- Any other relevant figures that help to understand the performance of the proposed method
- Make sure that the figures are properly labeled and include legends, titles, and axis labels. There should be data points in the figures, and the figures should be easy to read and interpret.
3. Write a README.md file to explain how to run the experiment.
4. Run the experiment automatically, visualize the results, and save the results. Make sure the generated figure files are not empty.
5. Save the experiment execution process in log.txt.
6. Analyze and summarize the experiment results in results.md after all experiments (including all baselines) have been successfully completed.
- Tables should be generated for the main results, including but not limited to:
- Summary of the experimental setup (e.g., hyperparameters, dataset splits)
- Comparison of the proposed method with baseline methods
- Any other relevant tables that help to understand the performance of the proposed method
7. Finally, create a folder named ’results’ under this folder and move results.md, log.txt, and all figures generated by the experiment into ’results’.
IMPORTANT:
- Do not use synthetic results or generate any fake data. The results should be based on real experiments. If the experimental dataset is too large, please use a smaller subset of the dataset for testing purposes.
- If the experiment is about large language models (LLMs), please confirm the models used are LLMs and not simple multi-layer neural networks such as LSTM.
- You can download the datasets from Hugging Face datasets or other open-sourced datasets. Please do not use any closed-sourced datasets.
- If the experiment requires using open-sourced large language models (LLMs) such as Meta Llama-3.1-8B, Llama-3.2-1B, Mistral Ministral-8B or Qwen Qwen3-0.6B and multimodal LLMs like Qwen2-VL-3B, please download the models from the Hugging Face model hub. Due to limited computing resources, please use the models no larger than 8B.
-If the experiment requires using closed-sourced large language models (LLMs) such as OpenAI’s GPT-4o-mini, o4-mini or Claude-3.7-sonnet, please directly use the API provided by the model provider. API keys have already been provided in the environment variables. Please use the API key to access the model and run the experiment.
-Only save essential files that are necessary for the experiment. Do not create or save any unnecessary files, temporary files, or intermediate files. Keep the output minimal and focused on the core experiment requirements.
-No need to generate any paper or academic document. Focus only on the experimental implementation and results.
-Please remove checkpoints or datasets larger than 1MB after the experiment is completed.
Remember to analyze and summarize the experiment results, create a folder named ’results’ under this folder and move results.md, log.txt, and all figures generated by the experiment into ’results’.
- Figures and tables generated by the experiment should be included in the results.md with clear explanations of what they show. Please make sure the generated figures are not repeated. Please do not use the same figures multiple times in results.md.
- Make sure paths to the figures in results.md are correct and point to the right locations.
- The results.md should also include a discussion of the results, including any insights gained from the experiment and how they relate to the hypothesis.
- The results.md should include a summary of the main findings and conclusions drawn from the experiment.
- The results.md should also include any limitations of the experiment and suggestions for future work.
Table 33: The Google Form interface used to collect review scores, following the same reviewer rubrics as MLR-Judge (LABEL:tab:end2end_eval_prompt).
Prompt You are an excellent machine learning researcher!
Given the task, research idea, literature review, proposal and experiment results, please write a paper for the machine learning project.
It should include the following sections:
1. Title and Abstract: A concise title and a brief abstract summarizing the research.
2. Introduction: An introduction to the problem, its significance, and the proposed solution.
3. Related Work: A review of existing literature and how your work fits into the current landscape.
4. Methodology: A detailed description of the proposed method, including any algorithms or models used.
5. Experiment Setup: A description of the experimental setup, including datasets, metrics, and evaluation methods.
6. Experiment Results: A presentation of the results obtained from the experiments, including tables and figures.
7. Analysis: An analysis of the results, discussing their implications and any limitations.
8. Conclusion: A summary of the findings and suggestions for future work.
9. References: A list of references cited in the paper.
Figures and tables should be included in the paper. You may refer to the figures and tables in the experiment results. If there is no image in the experiment results, please do not create or cite any fake figures. Please directly use the paths of the figures in the markdown file and do not use any placeholders.
When writing mathematical formulas, you should use LaTeX syntax. For inline formulas, use single dollar signs, for example: $x^{2}$ to represent x squared. For block equations, use double dollar signs at the beginning and end, for example: $x^{2}$.
If you need to write text in mathematical formulas, please avoid invalid escape characters.
The paper should be well-structured and clearly present the research findings.
Here is the task: {Task}
Here is the idea: {Idea}
Here is the literature review: {Related Work}
Here is a summary of the experiment results: {Experiment Results}
D.2 Prompts and Rubrics for MLR-Judge
MLR-Judge combines a set of carefully designed review rubrics with an LLM-based evaluator. It supports both an end-to-end evaluation pipeline—assessing an AI agent’s ability to complete a full research project—and a stepwise evaluation pipeline that independently evaluates performance on (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. In the following sections, we describe the specific rubrics used for each stage as well as for end-to-end evaluation. We then present the prompt used to operate the MLR-Judge.
D.2.1 Review Rubrics
We provide the rubrics used to guide evaluation for each stage:
  • The rubrics used to evaluate the idea generation quality are shown in LABEL:tab:review_code_prompt.
  • The rubrics used to evaluate the research proposal quality are shown in LABEL:tab:review_proposal_prompt.
  • The rubrics used to evaluate the experimentation quality are shown in LABEL:tab:review_code_prompt.
  • The rubrics used to evaluate the paper writing quality are shown in LABEL:tab:review_writing_prompt.
  • The rubric used for holistic paper evaluation is shown in LABEL:tab:end2end_eval_prompt.
Table 33: The Google Form interface used to collect review scores, following the same reviewer rubrics as MLR-Judge (LABEL:tab:end2end_eval_prompt).
1. CONSISTENCY (1-10)
How well does the idea align with the requirements of the task description?
9-10 - Excellent: The idea is perfectly aligned with the task description. It addresses all aspects of the task and is highly relevant.
7-8 - Good: The idea is mostly aligned with the task description. It addresses most aspects but may miss some minor details.
5-6 - Satisfactory: The idea is somewhat aligned with the task description. It addresses some aspects but misses several key points.
3-4 - Needs Improvement: The idea is poorly aligned with the task description. It addresses only a few aspects and misses many key points.
1-2 - Poor: The idea is not aligned with the task description. It does not address the task or is completely irrelevant.
2. CLARITY (1-10)
How clear and well-defined is the research idea?
9-10 - Excellent: The idea is crystal clear, perfectly defined, and immediately understandable. It is articulated concisely with no ambiguity.
7-8 - Good: The idea is mostly clear and well-articulated with only minor ambiguities. Minor refinements would make it even more precise.
5-6 - Satisfactory: The idea is partially clear but has some ambiguities. Some aspects need further elaboration for a complete understanding.
3-4 - Needs Improvement: The idea is unclear with significant ambiguities. Major clarification is needed to make it comprehensible.
1-2 - Poor: The idea is extremely unclear, vague, or ambiguous with no proper explanation. It is difficult to understand or interpret.
3. NOVELTY (1-10)
How original and innovative is the research idea?
9-10 - Excellent: The idea is highly original and innovative. It introduces a groundbreaking concept or approach that is significantly different from existing work.
7-8 - Good: The idea has notable originality and innovation. It offers fresh perspectives or new combinations of existing concepts.
5-6 - Satisfactory: The idea has some originality and innovation. It includes some novel aspects but also shares similarities with existing approaches.
3-4 - Needs Improvement: The idea has minimal originality. It largely resembles existing approaches with only slight variations.
1-2 - Poor: The idea lacks originality and innovation. It closely resembles common or existing concepts without any new insights.
4. FEASIBILITY (1-10)
How practical and implementable is the research idea?
9-10 - Excellent: The idea is highly practical and implementable with current resources, technology, and knowledge. Execution is straightforward.
7-8 - Good: The idea is largely feasible with existing technology and methods, though it may require moderate refinement or optimization.
5-6 - Satisfactory: The idea is somewhat feasible but has some implementation challenges. It may require considerable effort or resources to implement.
3-4 - Needs Improvement: The idea has significant implementation challenges. Major revisions would be needed to make it feasible.
1-2 - Poor: The idea is impractical or impossible to implement with current technology, knowledge, or constraints.
5. SIGNIFICANCE (1-10)
How important and impactful is the research idea?
9-10 - Excellent: The idea is highly significant and impactful. It addresses a critical problem and could lead to major advancements in the field.
7-8 - Good: The idea is significant and has clear impact potential. It addresses an important issue and could lead to meaningful contributions.
5-6 - Satisfactory: The idea is somewhat significant. It addresses a relevant problem but its impact may be moderate or limited to a specific area.
3-4 - Needs Improvement: The idea has limited significance. It addresses a minor issue or has a narrow scope with minimal impact.
1-2 - Poor: The idea has little to no significance. It does not address a meaningful problem or offer any clear benefits.
6. Overall Assessment (1-10)
How would you rate the research idea overall, considering all five dimensions above?
10 - Outstanding: The idea is exceptional in every respect, with no significant weaknesses.
8-9 - Excellent: The idea is very strong overall, with only minor weaknesses.
6-7 - Good: The idea is solid, with a good balance of strengths and weaknesses.
4-5 - Satisfactory: The idea is adequate but has notable weaknesses.
2-3 - Needs Improvement: The idea has significant weaknesses that limit its potential.
1 - Poor: The idea is fundamentally flawed across most or all dimensions.
When assigning the Overall Assessment score, consider not just the average of the five dimensions, but also:
- Whether any single weakness is critical enough to lower the overall potential.
- The overall coherence and integration of the idea.
- The likelihood of real-world impact if the idea were pursued.
- The degree to which the idea fulfills the task description as a whole.
- Any unique strengths or fatal flaws that are not fully captured by the individual dimensions.